# TaP: A Taxonomy-Guided Framework for Automated and Scalable Preference Data Generation

### v3

@@ 6.3 Evaluation
To evaluate the performance of trained LLMs, we employ MT-Bench-zh [^10] https://github.com/HIT-SCIR/huozi/tree/main/data/mt-bench-zh as the validation set, a Chinese translation of MT-Bench (DBLP:conf/nips/ZhengC00WZL0LXZ23). For testing, we adopt AlignBench (DBLP:conf/acl/LiuLWHFWCKXTZ0G24). Both benchmarks rely on high-performing LLMs, such as GPT-4o, to evaluate model responses. Accordingly, we report scores generated by both GPT-4o and DeepSeek-V3 to provide a comprehensive evaluation. Further details regarding the evaluation setup are provided in Appendix D.2.
@@ 7 Experimental Results
The experimental results presented in Table 4 suggest that despite the relatively small volume of TaP-SFT (GPT-4), LLMs trained on this dataset outperform those trained on other open-source datasets. Notably, LLMs trained on TaP-SFT (GPT-4) consistently surpass those trained on BELLE-SFT, even though BELLE-SFT contains over 180 times the number of samples as TaP-SFT (GPT-4). This finding indicates that, beyond dataset size, factors such as data quality, diversity, and coverage play a critical role in model performance. These results align with the findings of LIMA (DBLP:conf/nips/ZhouLX0SMMEYYZG23), which demonstrate that LLMs trained on just 1,000 carefully curated samples can achieve remarkably strong performance. However, LIMA’s heavy reliance on human curation poses challenges for scalability. In contrast, TaP necessitates human involvement only in designing the taxonomy, [^11] Instead of constructing the taxonomy manually, it can alternatively be constructed either by LLMs alone or through collaboration between humans and LLMs. making it significantly more scalable.
The capabilities of pre-trained LLMs serve as a critical foundation for the performance of models subsequently trained through supervised fine-tuning or preference fine-tuning.
The experimental results presented in Tables 4, 5, and 6 indicate that LLMs initialized from Qwen2.5-3B usually outperform those initialized from Llama-3.1-8B and Gemma-2-9B, despite the latter two LLMs having nearly three times the number of parameters as Qwen2.5-3B, even when trained on the same dataset. We hypothesize that this performance gap arises from the limited volume of Chinese corpora used in the pre-training of Llama-3.1-8B and Gemma-2-9B, which constrains their ability to generate high-quality responses in Chinese after undergoing supervised fine-tuning or preference fine-tuning. In contrast, the Qwen2.5 series of LLMs have been extensively trained on Chinese corpora, thereby enhancing their capacity for instruction following and alignment with human preferences and values after fine-tuning.
The effectiveness of applying supervised fine-tuning and preference fine-tuning using the same training dataset varies across different LLMs.
The experimental results presented in Tables 4, 5, and 6 indicate that the optimal open-source datasets for fine-tuning vary across LLMs. For instance, as shown in Table  4, the open-source fine-tuning dataset that yields the best performance for Qwen2.5-7B on MT-Bench-zh, when GPT-4o serves as the judge, is MOSS-SFT. However, for Llama-3.1-8B, the open-source dataset that results in the best performance under the same evaluation conditions is Infinity-Instruct. These findings align with the hypothesis proposed by zhang2025bestinstructiontuningdatafit, which posits that supervised fine-tuning datasets that are more closely aligned with the distribution of pre-trained LLMs tend to be more effective. Nevertheless, LLMs trained on data constructed by TaP consistently achieve strong performance and generally outperform those trained on other open-source datasets. This demonstrates that data constructed by TaP can be effectively employed for both supervised fine-tuning and preference fine-tuning across various LLMs.
@@ 8 Discussion
Although LLMs trained on the synthetic dataset constructed by TaP consistently outperform those trained on other open-source datasets, the underlying mechanisms responsible for this strong performance remain insufficiently understood. More broadly, while LLMs trained on different datasets can exhibit substantial performance differences, a comprehensive theoretical explanation of why training on one dataset yields superior performance compared to another remains an open problem, and a complete theoretical account of this phenomenon is not yet available. Nevertheless, we offer our current understanding as follows.
Recent studies have shown that LLMs benefit substantially from training data that are both diverse and knowledge-rich (lozhkov2024fineweb-edu; DBLP:journals/corr/abs-2309-05463; DBLP:conf/iclr/ZhangZZCWZBQ0FY25). We hypothesize that the superior performance of LLMs trained on the TaP-constructed synthetic dataset, relative to those trained on open-source datasets, can be attributed to several desirable properties of the samples it contains: (1) they are knowledge-rich, (2) they exhibit high information density, and (3) they span a broad range of domains and tasks. Figure 3 in Appendix B provides an illustrative example highlighting these properties. These desirable characteristics arise primarily from the taxonomy employed in the TaP framework to guide synthetic data generation. Specifically, the taxonomy is derived from the Undergraduate Program Catalog of Regular Higher Education Institutions of China, which constitutes a comprehensive and well-structured repository of foundational knowledge across disciplines. As a result, the synthetic samples generated by TaP capture a wide spectrum of essential knowledge, thereby making them particularly effective for training LLMs.
@@ 9 Conclusion
In this paper, we propose the TaP framework to facilitate automated and scalable construction of preference datasets across various languages. Grounded in a structured taxonomy, TaP offers fine-grained control over dataset composition, ensuring both diversity and comprehensive coverage. To evaluate the effectiveness of TaP and address the scarcity of preference datasets in Chinese, we apply TaP to construct preference datasets in Chinese, using a taxonomy derived from the Undergraduate Program Catalog of Regular Higher Education Institutions of China. We conduct supervised fine-tuning and preference fine-tuning on LLMs ranging from 3B to 14B parameters across three model families, utilizing the TaP-generated datasets. Experimental results show that LLMs trained on TaP-generated datasets outperform those trained on open-source datasets.
@@ Appendix A Alternative Approaches for Constructing the Taxonomy
While comprehensively enumerating and categorizing all possible use cases of LLMs is inherently challenging, a practical strategy is to collect user–model interactions and analyze these conversations to systematically identify and categorize use cases (DBLP:conf/iclr/Zhao0HC0D24; DBLP:conf/iclr/ZhengC0LZW00LXG24). For low-resource languages with limited resources, a feasible approach would be to initially deploy a LLM trained on all available data for these languages, potentially supplemented with data translated from high-resource languages. Following deployment, user–model conversations can be continuously collected and analyzed to construct the taxonomy grounded in real-world use cases. Additionally, both the taxonomy and the LLMs themselves can be incrementally updated as sufficient new user-model conversations become available, thereby improving model performance over time.
@@ Appendix B An Illustrative Example
Appendix C Dataset Construction
@@ D.1 Preference Fine-tuning
For PPO training, we first train the reward model to score the responses generated by the policy. Following common practice, we initialize the reward models from pre-trained LLMs (DBLP:journals/corr/abs-2311-04919; DBLP:conf/icml/CuiY0YH0NXXL0024; DBLP:conf/iclr/KirkMNLHGR24). Specifically, the reward models are initialized from the same pre-trained LLMs used to initialize the LLMs for supervised fine-tuning, with the exception that the reward model for the policy initialized from the checkpoints derived from supervised fine-tuning of Qwen2.5-3B is initialized from Qwen2.5-14B. Additionally, we constrain the rewards produced by the reward models to a predefined range $\left[-R,R\right]$ using an auxiliary loss function to stabilize PPO training, where $R$ is a positive constant. Given a prompt $x$, a preferred response $y_{w}$ and less preferred response $y_{l}$, the reward model is parameterized by $\phi$, with $\mathcal{D}$ denoting the preference dataset used for training. The auxiliary loss function is formulated as follows:
@@
For DPO training, we follow a similar strategy to PPO. Specifically, we train the policy for two epochs when using our constructed dataset and for three epochs when using open-source datasets. The learning rate follows the cosine schedule with a warm-up phase of 50 step and a peak learning rate of 2e-7. Following DBLP:conf/nips/IvisonW0WP0S0H24, we set $\beta$ to 0.01 across all experiments.
@@ Appendix E Experimental Results
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 6.29 3.36 4.83 6.39 2.98 4.68
COIG (DBLP:journals/corr/abs-2304-07987) 4.18 2.53 3.35 4.56 2.31 3.44
MOSS-SFT (MIR-2023-12-294) 6.31 3.00 4.66 6.01 2.74 4.38
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 5.99 2.95 4.47 6.01 2.69 4.35
@@
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 6.64 3.40 5.02 6.54 3.24 4.89
COIG (DBLP:journals/corr/abs-2304-07987) 4.40 2.55 3.48 4.86 2.46 3.66
MOSS-SFT (MIR-2023-12-294) 6.94 3.18 5.06 6.95 2.85 4.90
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 6.50 2.96 4.73 6.49 2.54 4.51
@@
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 5.46 2.95 4.21 5.66 2.90 4.28
COIG (DBLP:journals/corr/abs-2304-07987) 3.39 1.98 2.68 3.63 1.86 2.74
MOSS-SFT (MIR-2023-12-294) 5.28 3.15 4.21 5.35 2.76 4.06
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 4.36 2.26 3.32 4.66 2.03 3.34
@@
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 5.56 3.21 4.39 5.51 2.66 4.09
COIG (DBLP:journals/corr/abs-2304-07987) 3.35 2.03 2.69 3.49 1.81 2.65
MOSS-SFT (MIR-2023-12-294) 5.31 2.98 4.14 5.44 3.03 4.23
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 4.26 2.46 3.36 4.19 2.08 3.13
@@
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 7.14 3.75 5.44 7.01 3.71 5.36
COIG (DBLP:journals/corr/abs-2304-07987) 4.64 2.63 3.63 4.69 2.43 3.56
MOSS-SFT (MIR-2023-12-294) 6.99 3.53 5.26 7.00 3.08 5.04
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 6.75 3.09 4.92 6.74 2.83 4.78
@@
Qwen2.5-3B-SFT-Open Huozi-RLHF (huozi) 6.71 3.35 5.03 6.79 3.09 4.94
@@
Qwen2.5-3B-SFT-TaP Huozi-RLHF (huozi) 7.46 3.98 5.72 7.18 3.31 5.24
@@
Qwen2.5-7B-SFT-Open Huozi-RLHF (huozi) 7.09 3.41 5.25 6.96 3.00 4.98
@@
Qwen2.5-7B-SFT-TaP Huozi-RLHF (huozi) 8.10 4.10 6.10 7.99 3.71 5.85
@@
Llama-3.1-8B-SFT-Open Huozi-RLHF (huozi) 6.34 3.36 4.85 6.38 2.78 4.58
@@
Llama-3.1-8B-SFT-TaP Huozi-RLHF (huozi) 6.81 3.56 5.19 6.59 3.09 4.84
@@
Gemma-2-9B-SFT-Open Huozi-RLHF (huozi) 6.21 3.33 4.77 6.14 3.10 4.62
@@
Gemma-2-9B-SFT-TaP Huozi-RLHF (huozi) 6.95 3.84 5.39 6.81 3.21 5.01
@@
Qwen2.5-14B-SFT-Open Huozi-RLHF (huozi) 7.16 3.66 5.41 7.50 3.46 5.48
@@
Qwen2.5-14B-SFT-TaP Huozi-RLHF (huozi) 8.21 4.29 6.25 8.30 3.93 6.11
@@
Qwen2.5-3B-SFT-Open Huozi-RLHF (huozi) 6.86 3.64 5.25 6.91 3.49 5.20
@@
Qwen2.5-3B-SFT-TaP Huozi-RLHF (huozi) 7.51 4.21 5.86 7.65 3.86 5.76
@@
Qwen2.5-7B-SFT-Open Huozi-RLHF (huozi) 7.05 3.36 5.21 7.21 3.18 5.19
@@
Qwen2.5-7B-SFT-TaP Huozi-RLHF (huozi) 8.15 4.35 6.25 8.35 4.03 6.19
@@
Llama-3.1-8B-SFT-Open Huozi-RLHF (huozi) 6.34 3.53 4.93 6.46 3.33 4.89
@@
Llama-3.1-8B-SFT-TaP Huozi-RLHF (huozi) 6.61 3.89 5.25 6.85 3.24 5.04
@@
Gemma-2-9B-SFT-Open Huozi-RLHF (huozi) 6.41 3.33 4.87 6.49 3.21 4.85
@@
Gemma-2-9B-SFT-TaP Huozi-RLHF (huozi) 6.85 3.69 5.27 7.05 3.16 5.11
@@
Qwen2.5-14B-SFT-Open Huozi-RLHF (huozi) 7.59 3.98 5.78 7.80 3.76 5.78
@@
Qwen2.5-14B-SFT-TaP Huozi-RLHF (huozi) 8.29 4.54 6.41 8.60 4.43 6.51
@@
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 4.67 3.65 4.16 4.86 4.61 4.84 4.77 5.42 4.69 4.87 4.51
COIG (DBLP:journals/corr/abs-2304-07987) 2.96 2.49 2.73 4.02 4.44 3.68 3.84 4.61 4.16 4.12 3.42
MOSS-SFT (MIR-2023-12-294) 4.30 3.34 3.82 4.96 4.60 4.99 4.99 5.11 5.09 4.96 4.39
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 3.79 3.42 3.61 4.15 4.28 4.51 4.57 4.21 4.56 4.38 3.99
@@
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 5.49 4.30 4.90 5.50 5.21 5.44 5.24 5.13 5.34 5.31 5.10
COIG (DBLP:journals/corr/abs-2304-07987) 3.04 2.61 2.82 4.65 4.68 4.10 4.01 4.55 4.45 4.41 3.62
MOSS-SFT (MIR-2023-12-294) 5.10 4.32 4.71 5.46 4.78 4.87 5.35 5.68 5.41 5.26 4.98
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 5.26 4.17 4.72 5.10 5.35 5.07 5.15 5.13 4.96 5.13 4.92
@@
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 2.50 3.04 2.77 4.12 3.84 4.09 4.67 4.37 4.59 4.28 3.53
COIG (DBLP:journals/corr/abs-2304-07987) 1.73 1.85 1.79 2.96 2.53 2.90 2.79 3.32 3.37 2.98 2.38
MOSS-SFT (MIR-2023-12-294) 2.63 3.05 2.84 4.23 3.66 4.34 4.62 4.97 4.66 4.41 3.63
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 2.30 2.78 2.54 3.45 3.28 3.41 3.49 3.95 3.86 3.57 3.06
@@
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 3.18 3.16 3.17 4.32 3.36 4.18 4.40 4.47 4.59 4.22 3.70
COIG (DBLP:journals/corr/abs-2304-07987) 1.47 1.92 1.70 2.94 2.38 2.40 2.56 3.26 3.30 2.81 2.25
MOSS-SFT (MIR-2023-12-294) 2.72 2.96 2.84 4.00 3.38 3.89 4.55 5.16 4.30 4.21 3.53
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 2.55 2.83 2.69 3.25 3.03 3.76 3.53 3.87 4.05 3.58 3.14
@@
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 6.14 5.13 5.64 5.94 5.53 6.12 5.37 5.84 5.59 5.73 5.68
COIG (DBLP:journals/corr/abs-2304-07987) 3.21 3.39 3.30 4.92 5.04 4.94 4.43 5.18 4.55 4.84 4.07
MOSS-SFT (MIR-2023-12-294) 5.30 5.20 5.25 5.57 5.05 5.75 5.75 5.74 5.66 5.59 5.42
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 5.55 4.73 5.14 5.44 5.36 5.52 5.25 4.97 5.04 5.27 5.20
@@
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 4.53 3.45 3.99 4.54 4.21 4.70 4.77 5.24 5.03 4.75 4.37
COIG (DBLP:journals/corr/abs-2304-07987) 2.87 2.80 2.84 3.82 3.84 3.76 3.65 4.79 4.30 4.03 3.43
MOSS-SFT (MIR-2023-12-294) 4.24 3.22 3.73 4.78 4.24 4.66 5.00 5.13 5.09 4.82 4.27
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 3.81 3.49 3.65 3.99 4.12 4.43 4.41 4.42 4.51 4.31 3.98

### v4

@@ 6.3 Evaluation
To evaluate the performance of trained LLMs, we employ MT-Bench-zh [^10] https://github.com/HIT-SCIR/huozi/tree/main/data/mt-bench-zh as the validation set, a Chinese translation of MT-Bench (Zheng et al., 2023). For testing, we adopt AlignBench (Liu et al., 2024). Both benchmarks rely on strong LLMs (e.g., GPT-4o) to evaluate model responses. Accordingly, we report scores from both GPT-4o and DeepSeek-V3 for a comprehensive evaluation. Further details on the evaluation setup are provided in Appendix D.2.
@@ 7 Experimental Results
[…]
The capabilities of pre-trained LLMs provide a critical foundation for the performance of models subsequently trained through supervised fine-tuning or preference fine-tuning.
The experimental results in Tables 4, 5, and 6 indicate that LLMs initialized from Qwen2.5-3B generally outperform those initialized from Llama-3.1-8B and Gemma-2-9B, even though the latter two have nearly three times as many parameters as Qwen2.5-3B when trained on the same dataset. We hypothesize that this gap arises from the limited volume of Chinese corpora used to pre-train Llama-3.1-8B and Gemma-2-9B, which constrains their ability to generate high-quality Chinese responses after supervised fine-tuning or preference fine-tuning. In contrast, the Qwen2.5 series of LLMs have been extensively trained on Chinese corpora, enhancing their capacity to follow instructions and better aligning with human preferences and values.
The effectiveness of applying supervised fine-tuning and preference fine-tuning using the same training dataset varies across LLMs.
The experimental results presented in Tables 4, 5, and 6 indicate that the optimal open-source datasets for fine-tuning vary across LLMs. For instance, as shown in Table 4, when GPT-4o serves as the judge, the open-source fine-tuning dataset that yields the best performance for Qwen2.5-7B on MT-Bench-zh is MOSS-SFT. However, for Llama-3.1-8B, the open-source dataset that yields the best performance under the same evaluation conditions is Infinity-Instruct. These findings align with the claim proposed by Zhang et al. (2025b), which posits that supervised fine-tuning datasets that are more closely aligned with the distribution of pre-trained LLMs tend to be more effective. Nevertheless, LLMs trained on data constructed by TaP consistently achieve strong performance and generally outperform those trained on other open-source datasets. These results demonstrates that data constructed by TaP can be effectively employed for both supervised fine-tuning and preference fine-tuning across a range of LLMs.
@@ 8 Discussion
Although LLMs trained on the synthetic dataset constructed by TaP consistently outperform those trained on other open-source datasets, the mechanisms behind this strong performance remain insufficiently understood. More broadly, while LLMs trained on different datasets can exhibit substantial performance differences, a comprehensive theory explaining why training on one dataset yields better performance than another remains an open problem, and a complete theoretical account of this phenomenon is still lacking. Nevertheless, we offer our current understanding as follows.
Recent studies have shown that LLMs benefit substantially from training data that are both diverse and knowledge-rich (Li et al., 2023; Zhang et al., 2025a). Similarly, Lozhkov et al. (2024); Allal et al. (2025) construct large-scale corpora by selectively retaining samples with high educational value. We hypothesize that the superior performance of LLMs trained on the TaP-constructed synthetic dataset, relative to those trained on open-source datasets, can be attributed to several desirable properties of the samples it contains: (1) they are knowledge-rich, (2) they exhibit high information density, and (3) they span a broad range of domains and tasks. Figure 3 in Appendix B provides an illustrative example highlighting these properties. These desirable characteristics arise primarily from the taxonomy employed in the TaP framework to guide synthetic data generation. Specifically, the taxonomy is derived from the Undergraduate Program Catalog of Regular Higher Education Institutions of China, which constitutes a comprehensive and well-structured repository of foundational knowledge across disciplines. As a result, the synthetic samples generated by TaP capture a wide spectrum of essential knowledge, thereby making them particularly effective for training LLMs.
@@ 9 Conclusion
In this paper, we propose the TaP framework to enable the automated, scalable construction of preference datasets across languages. Grounded in a structured taxonomy, TaP offers fine-grained control over dataset composition, ensuring both diversity and broad coverage. To evaluate TaP and address the scarcity of Chinese preference data, we apply TaP to construct Chinese preference datasets using a taxonomy derived from the Undergraduate Program Catalog of Regular Higher Education Institutions of China. We perform supervised fine-tuning and preference fine-tuning on LLMs ranging from 3B to 14B parameters across three model families using the TaP-generated datasets. Experimental results show that LLMs trained on TaP-generated datasets outperform those trained on open-source datasets.
Limitations
Previous studies have highlighted the risk of model collapse when training LLMs on synthetic data (Gerstgrasser et al., 2024; Kazdan et al., 2025; Shumailov et al., 2024). As the TaP framework can be employed to generate synthetic data, LLMs trained on TaP-constructed datasets may also be susceptible to model collapse; this is a limitation of TaP. Nevertheless, our work focuses on automated and scalable preference dataset construction rather than mitigating the risks associated with synthetic data. Substantial research has been devoted to preventing model collapse caused by synthetic data, such as mixing real and synthetic data (Seddik et al., 2024), selecting samples based on entropy metrics (Shi et al., 2025), and proposing new training objectives to counteract overconfidence (Shabgahi et al., 2025). Addressing these risks is beyond the scope of this study, and we leave it for future work.
@@ Appendix A Alternative Approaches for Constructing the Taxonomy
While comprehensively enumerating and categorizing all possible use cases of LLMs is inherently challenging, a practical strategy is to collect user–model interactions and analyze these conversations to systematically identify and categorize use cases (Zhao et al., 2024; Zheng et al., 2024). For low-resource languages with limited resources, a feasible approach would be to initially deploy a LLM trained on all available data for these languages, potentially supplemented with data translated from high-resource languages. Following deployment, user–model conversations can be continuously collected and analyzed to construct the taxonomy grounded in real-world use cases. Additionally, both the taxonomy and the LLMs themselves can be incrementally updated as sufficient new user-model conversations become available, thereby improving model performance over time.
@@ Appendix B An Illustrative Example
Appendix C Preference Dataset Curation
@@ D.1 Preference Fine-tuning
For PPO training, we first train the reward model to score the responses generated by the policy. Following common practice, we initialize the reward models from pre-trained LLMs (Gooding and Mansoor, 2023; Cui et al., 2024; Kirk et al., 2024). Specifically, the reward models are initialized from the same pre-trained LLMs used to initialize the LLMs for supervised fine-tuning, with the exception that the reward model for the policy initialized from the checkpoints derived from supervised fine-tuning of Qwen2.5-3B is initialized from Qwen2.5-14B. Additionally, we constrain the rewards produced by the reward models to a predefined range $\left[-R,R\right]$ using an auxiliary loss function to stabilize PPO training, where $R$ is a positive constant. Given a prompt $x$, a preferred response $y_{w}$ and less preferred response $y_{l}$, the reward model is parameterized by $\phi$, with $\mathcal{D}$ denoting the preference dataset used for training. The auxiliary loss function is formulated as follows:
@@
For DPO training, we follow a similar strategy to PPO. Specifically, we train the policy for two epochs when using our constructed dataset and for three epochs when using open-source datasets. The learning rate follows the cosine schedule with a warm-up phase of 50 step and a peak learning rate of 2e-7. Following Ivison et al. (2024), we set $\beta$ to 0.01 across all experiments.
@@ Appendix E Experimental Results
Alpaca-GPT-4-ZH (Peng et al., 2023) 6.29 3.36 4.83 6.39 2.98 4.68
COIG (Zhang et al., 2023a) 4.18 2.53 3.35 4.56 2.31 3.44
MOSS-SFT (Sun et al., 2024) 6.31 3.00 4.66 6.01 2.74 4.38
COIG-CQIA (Bai et al., 2024) 5.99 2.95 4.47 6.01 2.69 4.35
@@
Alpaca-GPT-4-ZH (Peng et al., 2023) 6.64 3.40 5.02 6.54 3.24 4.89
COIG (Zhang et al., 2023a) 4.40 2.55 3.48 4.86 2.46 3.66
MOSS-SFT (Sun et al., 2024) 6.94 3.18 5.06 6.95 2.85 4.90
COIG-CQIA (Bai et al., 2024) 6.50 2.96 4.73 6.49 2.54 4.51
@@
Alpaca-GPT-4-ZH (Peng et al., 2023) 5.46 2.95 4.21 5.66 2.90 4.28
COIG (Zhang et al., 2023a) 3.39 1.98 2.68 3.63 1.86 2.74
MOSS-SFT (Sun et al., 2024) 5.28 3.15 4.21 5.35 2.76 4.06
COIG-CQIA (Bai et al., 2024) 4.36 2.26 3.32 4.66 2.03 3.34
@@
Alpaca-GPT-4-ZH (Peng et al., 2023) 5.56 3.21 4.39 5.51 2.66 4.09
COIG (Zhang et al., 2023a) 3.35 2.03 2.69 3.49 1.81 2.65
MOSS-SFT (Sun et al., 2024) 5.31 2.98 4.14 5.44 3.03 4.23
COIG-CQIA (Bai et al., 2024) 4.26 2.46 3.36 4.19 2.08 3.13
@@
Alpaca-GPT-4-ZH (Peng et al., 2023) 7.14 3.75 5.44 7.01 3.71 5.36
COIG (Zhang et al., 2023a) 4.64 2.63 3.63 4.69 2.43 3.56
MOSS-SFT (Sun et al., 2024) 6.99 3.53 5.26 7.00 3.08 5.04
COIG-CQIA (Bai et al., 2024) 6.75 3.09 4.92 6.74 2.83 4.78
@@
Qwen2.5-3B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.71 3.35 5.03 6.79 3.09 4.94
@@
Qwen2.5-3B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 7.46 3.98 5.72 7.18 3.31 5.24
@@
Qwen2.5-7B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 7.09 3.41 5.25 6.96 3.00 4.98
@@
Qwen2.5-7B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 8.10 4.10 6.10 7.99 3.71 5.85
@@
Llama-3.1-8B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.34 3.36 4.85 6.38 2.78 4.58
@@
Llama-3.1-8B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.81 3.56 5.19 6.59 3.09 4.84
@@
Gemma-2-9B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.21 3.33 4.77 6.14 3.10 4.62
@@
Gemma-2-9B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.95 3.84 5.39 6.81 3.21 5.01
@@
Qwen2.5-14B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 7.16 3.66 5.41 7.50 3.46 5.48
@@
Qwen2.5-14B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 8.21 4.29 6.25 8.30 3.93 6.11
@@
Qwen2.5-3B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.86 3.64 5.25 6.91 3.49 5.20
@@
Qwen2.5-3B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 7.51 4.21 5.86 7.65 3.86 5.76
@@
Qwen2.5-7B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 7.05 3.36 5.21 7.21 3.18 5.19
@@
Qwen2.5-7B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 8.15 4.35 6.25 8.35 4.03 6.19
@@
Llama-3.1-8B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.34 3.53 4.93 6.46 3.33 4.89
@@
Llama-3.1-8B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.61 3.89 5.25 6.85 3.24 5.04
@@
Gemma-2-9B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.41 3.33 4.87 6.49 3.21 4.85
@@
Gemma-2-9B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.85 3.69 5.27 7.05 3.16 5.11
@@
Qwen2.5-14B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 7.59 3.98 5.78 7.80 3.76 5.78
@@
Qwen2.5-14B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 8.29 4.54 6.41 8.60 4.43 6.51
@@
Alpaca-GPT-4-ZH (Peng et al., 2023) 4.67 3.65 4.16 4.86 4.61 4.84 4.77 5.42 4.69 4.87 4.51
COIG (Zhang et al., 2023a) 2.96 2.49 2.73 4.02 4.44 3.68 3.84 4.61 4.16 4.12 3.42
MOSS-SFT (Sun et al., 2024) 4.30 3.34 3.82 4.96 4.60 4.99 4.99 5.11 5.09 4.96 4.39
COIG-CQIA (Bai et al., 2024) 3.79 3.42 3.61 4.15 4.28 4.51 4.57 4.21 4.56 4.38 3.99
@@
Alpaca-GPT-4-ZH (Peng et al., 2023) 5.49 4.30 4.90 5.50 5.21 5.44 5.24 5.13 5.34 5.31 5.10
COIG (Zhang et al., 2023a) 3.04 2.61 2.82 4.65 4.68 4.10 4.01 4.55 4.45 4.41 3.62
MOSS-SFT (Sun et al., 2024) 5.10 4.32 4.71 5.46 4.78 4.87 5.35 5.68 5.41 5.26 4.98
COIG-CQIA (Bai et al., 2024) 5.26 4.17 4.72 5.10 5.35 5.07 5.15 5.13 4.96 5.13 4.92
@@
Alpaca-GPT-4-ZH (Peng et al., 2023) 2.50 3.04 2.77 4.12 3.84 4.09 4.67 4.37 4.59 4.28 3.53
COIG (Zhang et al., 2023a) 1.73 1.85 1.79 2.96 2.53 2.90 2.79 3.32 3.37 2.98 2.38
MOSS-SFT (Sun et al., 2024) 2.63 3.05 2.84 4.23 3.66 4.34 4.62 4.97 4.66 4.41 3.63
COIG-CQIA (Bai et al., 2024) 2.30 2.78 2.54 3.45 3.28 3.41 3.49 3.95 3.86 3.57 3.06
@@
Alpaca-GPT-4-ZH (Peng et al., 2023) 3.18 3.16 3.17 4.32 3.36 4.18 4.40 4.47 4.59 4.22 3.70
COIG (Zhang et al., 2023a) 1.47 1.92 1.70 2.94 2.38 2.40 2.56 3.26 3.30 2.81 2.25
MOSS-SFT (Sun et al., 2024) 2.72 2.96 2.84 4.00 3.38 3.89 4.55 5.16 4.30 4.21 3.53
COIG-CQIA (Bai et al., 2024) 2.55 2.83 2.69 3.25 3.03 3.76 3.53 3.87 4.05 3.58 3.14
@@
Alpaca-GPT-4-ZH (Peng et al., 2023) 6.14 5.13 5.64 5.94 5.53 6.12 5.37 5.84 5.59 5.73 5.68
COIG (Zhang et al., 2023a) 3.21 3.39 3.30 4.92 5.04 4.94 4.43 5.18 4.55 4.84 4.07
MOSS-SFT (Sun et al., 2024) 5.30 5.20 5.25 5.57 5.05 5.75 5.75 5.74 5.66 5.59 5.42
COIG-CQIA (Bai et al., 2024) 5.55 4.73 5.14 5.44 5.36 5.52 5.25 4.97 5.04 5.27 5.20
@@
Alpaca-GPT-4-ZH (Peng et al., 2023) 4.53 3.45 3.99 4.54 4.21 4.70 4.77 5.24 5.03 4.75 4.37
COIG (Zhang et al., 2023a) 2.87 2.80 2.84 3.82 3.84 3.76 3.65 4.79 4.30 4.03 3.43
MOSS-SFT (Sun et al., 2024) 4.24 3.22 3.73 4.78 4.24 4.66 5.00 5.13 5.09 4.82 4.27
COIG-CQIA (Bai et al., 2024) 3.81 3.49 3.65 3.99 4.12 4.43 4.41 4.42 4.51 4.31 3.98